OpenStudy (kaiiskool):

Please help and no straight answers. 10x^3-25x over 5x Thanks for the help!

OpenStudy (abdullahm):

Hint: \(\Large\bf \frac{10x^3-25x}{5x} = \frac{10x^3}{5x} - \frac{25x}{5x}\) What is 10x^3/5x = ? What is 25x/5x = ?

OpenStudy (kaiiskool):

Im guessing that the answer is 2^3 for the first one and 5 for the second

OpenStudy (abdullahm):

The second part is correct For the first one, let's break it down more so you can get it: \(\Large\bf \frac{2 * 5 * x* x* x}{5* x}\) How many x's can we take out from the top and the bottom of the fraction?

OpenStudy (kaiiskool):

We can only take two out? I think this because There must be a remainder so we know that its unsolved

OpenStudy (abdullahm):

We can take only 1 x out because there is 1 x in the denominator and one x in the numerator.  Do you understand now?